OpenStudy (korosh23):

Find a formula for the inverse of the function. f(x) = e^5x − 7 f^-1(x) = ? Please wait until I post my own attempt.

OpenStudy (korosh23):

This is the question: \[f(x)=e^{5x-7}\] \[f^-1(x)=?\] My own attempt: \[\frac{ \log_{e (y) +7} }{5 } = x\]

OpenStudy (korosh23):

I know my solution is wrong, but I cannot understand where is my mistake?

OpenStudy (bobo-i-bo):

Is it \(e^{5x-7}\) or \(e^{5x}-7\)?

OpenStudy (korosh23):

the first one you said

OpenStudy (bobo-i-bo):

then it's because you've done stuff in the wrong order... working out please? :)

OpenStudy (korosh23):

what should I do then?

OpenStudy (bobo-i-bo):

please show me your working out

OpenStudy (korosh23):

y = e^5x-7 f^-1(y) = e^5x -7

OpenStudy (bobo-i-bo):

please show all the step by step working. That will show you where you went wrong :)